You are here: Foswiki>Commissioning Web>Logs>30Mar2012_Log (31 Mar 2012, JuanGuerra)Edit Attach

AO Commissioning on 30 March 2012

Observer(s):C. Arcidiacono, L. Busoni, A.Riccardi,F. Quiros JC. Guerra (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: J. Urban (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Close dome calibrations using the retroreflector.

The awake up of the left system. OK. Hexapod alignment using the hexapod spiral IDL procedure. Calibration of the Bin3, acquiring the pupils and intmat and building the Reconstructor. The reconstructor has been tested with the atmosphere disturbance giving problems that RIP the shell. It was identify an actuator(#20) which had a bad behaviour, this actuator was disabled from the list. The hexapod alignment took most of the daytime activities.

Investigating with Armando the problems during te SET of the left shell. We have decided that this could be due to a bad position of the actuators 20 and 583. We couldn't find why they jump on/off during the SET process. We carried out several test of REST/SET the shell and we haven't find any problem (before and after remove teh act #20).

Power on the W#1. Ok. Checking the temperature trend of the BCU47.

Details

8:00 Power on the system, OK.

08:30 Alignment the hexapod.

Running the procedure hexa_spiral to move the hexapod and to identidy where and which are the right spots. This procedure took most of the morning and the spots were found far from the original position

After a first cycle of spiral some reflexions were seen in the IRTC (wide FOV) and due to the long time taking for a cycle, it was decided to increase the delta of the movements. Using this method we could see some defocused image up of the IRTC (wide FOV). The procedure was stopped and a new spiral was done using a different center close to the position were we saw the spots. The spots were identified with a visible coma aberration. Once the spot is identified ( right) is centered in the center of the detector which correspods with the center of the narrow field. And the spot is centersd in the pyramid.

  • IRTC_spots_coma:
    comaIRTC.png

The IRTC FOV is changed to narrow and the coma aberration is compensated manually using the PSF and the use of the stages to moved to center the spot in the CCD47 using the X and Y offset from the W adsec GUI. The spot in the IRTC is focused using the stage. Focus Z=39.8mm

A board setup is saved to keep the actual W# cofiguration. Board setup saved: telescope-align5

Hexapod position:

x(mm) y(mm) z(mm) Rx(arcsec) Ry(arcsec) Rz(arcsec)
4.6 1.5 4.58 317 612 0

BIN 3 calibration

The reconstuctor for the Bin 3 was built and it wasn't tested due to problems with the secondary. The close loop with the disturbace load and running bin3 the secondary RIP, it has been tested many times and the cause of all RIPs were some bad behaviour of the some actuators actuators. The actuator were identified and one of them was removed from the configurations files. Actuator #20. After remove the actuator we tried to close the loop again and the system didn't response well. This behavious has been investigated off-line. An optimization of the BIN3 at different flux is programed for next day.

  • IRTC_10modes:
    IRTC_10modes.png
Topic revision: r4 - 31 Mar 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback